# Robust Bayesian Inference for Berkson and Classical Measurement Error Models

### v2

@@ B.2 Proofs of theoretical results for the Classical error model
$\operatorname{\operatorname{MMD}}_{k^{2}_{X}}(\operatorname{\mathbb{P}}_{X}^{0},\operatorname{\mathbb{F}})=0\quad\text{if and only if}\operatorname{\mathbb{P}}_{X}^{0}=\operatorname{\mathbb{F}}.$
Appendix C Experimental setup & additional results
In this section, we provide details of our experimental setup in Section 5. For all applications of the Romust-MEM (MMD) method, we used automatic differentiation with the Adam optimizer (Kingma and Ba,, 2014) provided by JAX, to minimize the MMD objective which is approximated via a U-statistic as in Gretton et al., (2012) using independent samples. In particular, for i.i.d samples $\{x_{i}\}_{i=1}^{n}\operatorname{\overset{iid}{\sim}}\operatorname{\mathbb{P}}$ and $\{y_{i}\}_{i=1}^{m}\operatorname{\overset{iid}{\sim}}\operatorname{\mathbb{Q}}$ the squared MMD between $\operatorname{\mathbb{P}},\operatorname{\mathbb{Q}}$ can be approximated as:
$\widehat{\operatorname{\operatorname{MMD}}}^{2}_{k}(\operatorname{\mathbb{P}},\operatorname{\mathbb{Q}})=\frac{1}{n(n-1)}\sum_{i=1}^{n}\sum_{j\neq i}^{n}k(x_{i},x_{j})+\frac{1}{m(m-1)}\sum_{i=1}^{m}\sum_{j\neq i}^{m}k(y_{i},y_{j})-\frac{2}{nm}\sum_{i=1}^{n}\sum_{j=1}^{m}k(x_{i},y_{j}).$
@@ Appendix C Experimental setup & additional results
In practice, we take $n=m$ to be equal to the number of observations. To initialize the optimization several different approaches can be taken. In the nonlinear regression simulation, we sample uniformly at random from $[-1,4]$ at each bootstrap iteration for each regression coefficient. One could also initialize from the nonlinear Least Squares (NLS) estimate however adding randomness to the initialization procedure can aid numerical stability and increase posterior uncertainty leading to better coverage probabilities. In the real-world California Test Scores study we combine the above methods by first finding the NLS estimates and then sampling uniformly from intervals centered around these estimates. Initialization for the Mental Health study which has a more complicated model and a higher number of parameters is discussed in Section C.3.
@@
We further used the RBF kernel for both $k_{X}$ and $k_{Y}$ throughout our experiments. For the nonlinear regression simulation the length scales for $k_{X}$ and $k_{Y}$ were set to 100 in the Classical error case, whereas the median heuristic (Gretton et al.,, 2012) was used in the Berkson case as well as the California Test Scores study. For the Mental Health study the length scales for $k_{X}$ and $k_{Y}$ were set to 10 and 100 repsectively. These values were chosen to ensure the loss and gradient do not vanish during optimisation.
C.1 Additional results for the Linear Regression case
@@
We present the results of the Linear Regression simulation from section 5.1 in the well-specified case, i.e. when $\operatorname{\mathbb{F}}_{N}=\operatorname{\mathbb{F}}_{N}^{0}:=N(0,\sigma_{\nu}^{0})$ and $\operatorname{\mathbb{F}}_{X}=\operatorname{\mathbb{P}}_{X}^{0}:=N(0,1)$, for $\sigma_{\nu}^{0}\in\{10^{-6},0.5,1,2\}$. The linear regression fit is illustrated in Figure 4 for all methods. As expected, the SIMEX method performs better than in the misspecified case whereas all other methods perform similarly.
@@
Robust-MEM (TLS) $\theta_{1}$ 1.615 (0.129) 1.603 (0.166) 1.548 (0.273) 1.319 (0.573)
$\theta_{2}$ 0.013 (0.002) 0.013 (0.003) 0.013 (0.004) 0.012 (0.006)
@@
C.2 Additional results for the nonlinear Regression case with Berkson error
Similarly, we consider the example of nonlinear regression with Berkson ME from section 5.1 in the well-specified case where both the true and the prior ME distributions are Gaussian with increasing variance. Results are reported for 100 simulations in Figure 10.
@@ C.2 Additional results for the nonlinear Regression case with Berkson error
We observe very similar behaviour to the misspecified case. For ME variances of $\leq 0.25$, all methods exhibit comparable performance. However, as the variance escalates, the Robust-MEM (MMD) approach stands out as the most resilient to ME effects. This is also seen by observing the associated MSE and their variances in table 10.
@@
Robust-MEM (MMD) $\theta_{1}$ 0.139 (0.24) 0.127 (0.22) 0.13 (0.179) 0.244 (0.279)
$\theta_{2}$ 0.442 (0.346) 0.238 (0.278) 0.4 (0.613) 1.264 (1.223)
@@
C.3 Mental Health Study
In this section we give details of the model used in Harezlak et al., (2018) for the treatment effect estimation in the Mental Health study which we also adopt in Section 5.2.
In particular, for each observation $i=1,\dots,n_{placebo}$ in the placebo group we model
@@ C.3 Mental Health Study
We define the parameter of interest $\theta=(\beta,u_{0},u_{1})$ where $\beta=(\beta_{0},\beta_{1},\beta_{0}^{\text{drug}},\beta_{1}^{\text{drug}})$, $u_{0}=(u_{01},\dots,u_{0K})$ and $u_{1}=(u_{11},\dots,u_{1K})$. Let $n=n_{treatment}+n_{placebo}$ and define $Y=(y_{1},\dots,y_{n})$ and the design matrices
$x=$ $\begin{pmatrix}1&x_{1}&1-I_{1}&(1-I_{1})x_{1}\\ \vdots&\vdots&\vdots&\vdots\\ 1&x_{n}&1-I_{n}&(1-I_{n})x_{n}\end{pmatrix},\quad Z_{0}=\begin{pmatrix}I_{1}z_{1}(x_{1})&\dots&I_{1}z_{K}(x_{1})\\ \vdots&\ddots&\vdots\\ I_{n}z_{1}(x_{n})&\dots&I_{n}z_{K}(x_{n})\end{pmatrix}\quad\text{and}$
@@
For the implementation of the Robust-MEM (MMD) algorithm in this application we set $c=1$ and $T=100$. For the optimization step we used a learning rate of $\eta=0.001$ and we employ a random restart method for the initialisation of the optimization: 100 points are sampled uniformly at random and the MMD loss is calculated using the observed data. The point with the smallest loss is used as the initial point and the resulting parameter estimate with the smallest loss during optimisation is picked. The uniform intervals are set to $[0,2]$ for the regression parameters $\beta$ and $[-0.1,0.1]$ for the spline parameters $u_{0},u_{1}$. Posterior samples are obtained from 200 bootstrap iterations. Of course, when the optimisation objective is non-convex, there is usually a trade-off between coverage capability and MSE of the posterior estimates or intervals depending on the initialisation of the optimisation. A wider uniform interval usually leads to higher uncertainty in our posterior beliefs which can exhbit better coverage probability of estimated intervals. On the contrary, if one is interested in the MSE of the posterior estimations then one can perform the above procedure for more than 100 times to get a better initial point of the optimisation or use a narrower uniform distribution based on prior beliefs or historic data.

### v3

@@ D.3 Extension of theoretical results to the DP approximation
$\quad\geq\inf_{\hat{\theta}_{n}}\sup_{\begin{subarray}{c}{\operatorname{\mathbb{P}}^{0}_{\theta^{\star}}}:\\ \text{no ME}\end{subarray}}\operatorname{\mathbb{E}}_{(x_{i},y_{i})_{1:n}\sim\operatorname{\mathbb{P}}^{0}_{\theta^{\star}}}[\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}^{0}_{\theta^{\star}},\operatorname{\mathbb{P}}_{\hat{\theta}_{n}})]$
$\quad=\inf_{\hat{\theta}_{n}}\sup_{\begin{subarray}{c}{\operatorname{\mathbb{P}}^{0}_{\theta^{\star}}}\end{subarray}}\operatorname{\mathbb{E}}_{(x_{i},y_{i})_{1:n}\sim\operatorname{\mathbb{P}}^{0}_{\theta^{\star}}}[\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}^{0}_{\theta^{\star}},\operatorname{\mathbb{P}}_{\hat{\theta}_{n}})].$
We can now use Le Cam’s method. Let $\operatorname{\mathbb{P}}_{g(\theta,X)}:=\mathcal{N}(\theta,\sigma^{2})$ and let $\theta_{0}=D$ and $\theta_{1}=-D$ for some $D>0$. Denote the joint likelihood across $n$ observations $(x_{i},y_{i})_{1:n}$ as $\operatorname{\mathbb{P}}_{\theta}^{\times n}$. Then we have that:
$KL(\operatorname{\mathbb{P}}^{\times n}_{\theta_{0}}||\operatorname{\mathbb{P}}^{\times n}_{\theta_{0}})$ $=nKL(\operatorname{\mathbb{P}}_{X}^{0}\times\mathcal{N}(D,\sigma^{2})||\operatorname{\mathbb{P}}_{X}^{0}\times\mathcal{N}(-D,\sigma^{2}))$
$=nKL(\mathcal{N}(D,\sigma^{2})||\mathcal{N}(-D,\sigma^{2}))$
$=\frac{2nD^{2}}{\sigma^{2}}$
where the first equality follows from the chain rule for the KL divergence, the second equality follows from the fact that under this construction $X$ and $Y$ are independent and hence by log properties:
$KL(\operatorname{\mathbb{P}}_{X}^{0}\times\mathcal{N}(D,\sigma^{2})||\operatorname{\mathbb{P}}_{X}^{0}\times\mathcal{N}(-D,\sigma^{2}))$
$=KL(\operatorname{\mathbb{P}}_{X}^{0},\operatorname{\mathbb{P}}_{X}^{0})+KL(\mathcal{N}(D,\sigma^{2})||\mathcal{N}(-D,\sigma^{2}))$
$=KL(\mathcal{N}(D,\sigma^{2})||\mathcal{N}(-D,\sigma^{2}))$
and the third equality follows from the closed form expression of KL between Gaussian distributions. By Pinsker’s inequality (see e.g. Tsybakov,, 2008, Lemma 2.5), we hence have that:
$TV(\operatorname{\mathbb{P}}^{\times n}_{\theta_{0}},\operatorname{\mathbb{P}}^{\times n}_{\theta_{0}})\leq\sqrt{\frac{1}{2}KL(\operatorname{\mathbb{P}}^{\times n}_{\theta_{0}}||\operatorname{\mathbb{P}}^{\times n}_{\theta_{0}})}=\sqrt{n}\frac{D}{\sigma}.$
Now since $D$ is arbitrary, let us choose $D:=\frac{c\sigma}{\sqrt{n}}$ for some $c>0$. We then have $TV(\operatorname{\mathbb{P}}^{\times n}_{\theta_{0}},\operatorname{\mathbb{P}}^{\times n}_{\theta_{0}})\leq c$. Moreover, by Lemma D.9, we have that:
$\operatorname{\operatorname{MMD}}^{2}_{k}(\operatorname{\mathbb{P}}_{\theta_{0}},\operatorname{\mathbb{P}}_{\theta_{1}})$ $=\operatorname{\operatorname{MMD}}^{2}_{k}(\operatorname{\mathbb{P}}_{X}^{0}\times\mathcal{N}(D,\sigma^{2}),\operatorname{\mathbb{P}}_{X}^{0}\times\mathcal{N}(-D,\sigma^{2}))$
$=\operatorname{\mathbb{E}}_{x,x^{\prime}\sim\operatorname{\mathbb{P}}_{X}^{0}}[k_{X}(x,x^{\prime})\operatorname{\operatorname{MMD}}_{k_{Y}}^{2}(\mathcal{N}(D,\sigma^{2}),\mathcal{N}(-D,\sigma^{2}))]$
$=\operatorname{\mathbb{E}}_{x,x^{\prime}\sim\operatorname{\mathbb{P}}_{X}^{0}}[k_{X}(x,x^{\prime})]\operatorname{\operatorname{MMD}}_{k_{Y}}^{2}(\mathcal{N}(D,\sigma^{2}),\mathcal{N}(-D,\sigma^{2})).$
For the second term, we can use the closed form expression of the MMD for Gaussians with a Gaussian kernel as in the proofs of Corollaries 4.2.3 and 4.2.4 to obtain:
$\operatorname{\operatorname{MMD}}^{2}_{k_{Y}}(\mathcal{N}(D,\sigma^{2}),\mathcal{N}(-D,\sigma^{2}))=2\frac{l_{Y}}{\sqrt{l_{Y}^{2}+2\sigma^{2}}}\left(1-\exp\left(-\frac{2D^{2}}{l_{Y}^{2}+2\sigma^{2}}\right)\right).$
Let $t:=\frac{2D^{2}}{l_{Y}^{2}+2\sigma^{2}}=\frac{2c^{2}\sigma^{2}}{n(l_{Y}^{2}+2\sigma^{2})}$. Then if we choose $c>0$ such that $t\in[0,1]$ we have that $1-\exp(-t)\geq\frac{t}{2}$ and hence:
$\operatorname{\operatorname{MMD}}_{k_{Y}}(\mathcal{N}(D,\sigma^{2}),\mathcal{N}(-D,\sigma^{2}))\geq\sqrt{\frac{2l_{Y}}{(l_{Y}^{2}+2\sigma^{2})^{3/2}}}\frac{c\sigma}{\sqrt{n}}$
Now define $m_{X}:=\sqrt{\operatorname{\mathbb{E}}_{x,x^{\prime}\sim\operatorname{\mathbb{P}}_{X}^{0}}[k_{X}(x,x^{\prime})]}$ which is independent of $n$. We then have:
$\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}_{\theta_{0}},\operatorname{\mathbb{P}}_{\theta_{1}})\geq\frac{1}{\sqrt{n}}\left(m_{X}c\sigma\sqrt{\frac{2l_{Y}}{(l_{Y}^{2}+2\sigma^{2})^{3/2}}}\right).$
By Le Cam’s method (see Wainwright,, 2019, Lemma 15.9), we have that:
$\inf_{\hat{\theta}_{n}}\sup_{\begin{subarray}{c}{\operatorname{\mathbb{P}}^{0}_{\theta^{\star}}}\end{subarray}}\operatorname{\mathbb{E}}_{(x_{i},y_{i})_{1:n}\sim\operatorname{\mathbb{P}}^{0}_{\theta^{\star}}}[\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}^{0}_{\theta^{\star}},\operatorname{\mathbb{P}}_{\hat{\theta}_{n}})]$ $\geq\frac{MMD(\operatorname{\mathbb{P}}_{\theta_{0}},\operatorname{\mathbb{P}}_{\theta_{1}})}{4}(1-TV(\operatorname{\mathbb{P}}^{\times n}_{\theta_{0}},\operatorname{\mathbb{P}}^{\times n}_{\theta_{1}}))$
$\geq\frac{1}{\sqrt{n}}\left(m_{X}c\sigma\sqrt{\frac{2l_{Y}}{(l_{Y}^{2}+2\sigma^{2})^{3/2}}}\right)(1-c).$
Appendix E Experimental setup & additional results
In this section, we provide details of our experimental setup in Sections 5 and 6 as well as additional experimental results.
E.1 Experimental details
In the Linear Regression example of Section 5, for the Robust-MEM (TLS) algorithm We set $T=100$ and $c=1$ and vary $\sigma^{2}_{\nu}$ while keeping $\sigma_{\epsilon}$ fixed at 0.5 throughout the experiment. A pseudo-code of the Robust-MEM algorithm can be found in Alg. 3.
input : $(w_{i},y_{i})_{1:n}$, $T$, $B$, $c$, $(\operatorname{\mathbb{F}}_{w_{i}})_{1:n}$
for $j\leftarrow 1$ to $B$ do
 for $i\leftarrow 1$ to $n$ do
  Calculate $\theta^{(j)}=\theta^{\star}_{TLS}(\operatorname{\mathbb{P}}^{(1,j)},\dots,\operatorname{\mathbb{P}}^{(n,j)})$ as in (10)
return Posterior bootstrap sample $\theta^{(1:B)}$
Algorithm 3 Robust-MEM with the TLS
[…]
$\widehat{\operatorname{\operatorname{MMD}}}^{2}_{k}(\operatorname{\mathbb{P}},\operatorname{\mathbb{Q}})=\frac{1}{n(n-1)}\sum_{i=1}^{n}\sum_{j\neq i}^{n}k(x_{i},x_{j})+\frac{1}{m(m-1)}\sum_{i=1}^{m}\sum_{j\neq i}^{m}k(y_{i},y_{j})$
$\quad\quad-\frac{2}{nm}\sum_{i=1}^{n}\sum_{j=1}^{m}k(x_{i},y_{j}).$
@@ E.1 Experimental details
In practice, we take $n=m$ to be equal to the number of observations. To initialize the optimisation several different approaches can be taken. In the nonlinear regression simulation, we sample uniformly at random from [1,4][-1,4] at each bootstrap iteration for each regression coefficient. One could also initialize from the nonlinear Least Squares (NLS) estimate however adding randomness to the initialization procedure can aid numerical stability and increase posterior uncertainty leading to better coverage probabilities. In the real-world California Test Scores study we combine the above methods by first finding the NLS estimates and then sampling uniformly from intervals centered around these estimates. Initialization techniques for more complex models and a higher number of parameters, such as the Mental Health study, is discussed in Section E.5.
@@
Furthermore, the RBF kernel was used for both $k_{X}$ and $k_{Y}$ throughout our experiments. For the nonlinear regression simulation, the length scales for $k_{X}$ and $k_{Y}$ were set to 100 in the Classical error case, whereas the median heuristic Gretton et al., (2012) was used in the Berkson case as well as the California Test Scores study. For the Mental Health study, the length scales were 10 and 100 respectively. These values were chosen to ensure the loss and gradient do not vanish during optimisation. A pseudo-code of the Robust-MEM algorithm with the MMD-based loss function is outlined in 4.
input : $(w_{i},y_{i})_{1:n}$, $T$, $B$, $c$, $(\operatorname{\mathbb{F}}_{w_{i}})_{1:n}$
for $j\leftarrow 1$ to $B$ do
 for $i\leftarrow 1$ to $n$ do
  Obtain $\theta^{(j)}=\theta^{\ast}_{\text{MMD}}(\operatorname{\mathbb{P}}^{(1,j)},\dots,\operatorname{\mathbb{P}}^{(n,j)})$ as in 3.4.2 using numerical optimisation.
return Posterior bootstrap sample $\theta^{(1:B)}$
Algorithm 4 Robust-MEM with the MMD
E.2 Additional results for the Linear Regression case
We now consider the Linear regression simulation from Section 5. We first give additional results on the misspecified prior cases considered in the main text in Table 5. We observe that the Robust-MEM (TLS) method remains robust to ME even under a misspecified prior variance. In the absence of ME ($\sigma_{\nu}^{2}=1e-06$), OLS performs best whereas for small ME variance ($\sigma_{\nu}^{2}=0.5$) TLS achieves the smallest MSE with all methods performing well. The suggested method has the lowest MSE for higher values of the variance.
Table 5: MSE over 100 simulations with associated std. in the Linear Regression example with Classical ME (with a misspecified prior) for an increasing ME variance ($\sigma^{2}_{\nu}$).
Method ME variance
1e-06 0.5 1.0 2.0
OLS $\theta_{1}$ 0.063 (0.083) 1.632 (1.232) 16.945 (7.153) 103.85 (21.629)
$\theta_{2}$ 0.001 (0.001) 0.016 (0.012) 0.17 (0.071) 1.035 (0.209)
R-MEM $\theta_{1}$ 1.618 (0.128) 1.585 (0.161) 1.464 (0.222) 1.011 (0.335)
(TLS) $\theta_{2}$ 0.013 (0.002) 0.013 (0.003) 0.012 (0.004) 0.011 (0.008)
SIMEX $\theta_{1}$ 0.07 (0.073) 1.127 (0.681) 15.404 (4.712) 99.524 (16.695)
$\theta_{2}$ 0.001 (0.001) 0.011 (0.007) 0.154 (0.046) 0.993 (0.159)
TLS $\theta_{1}$ 0.118 (0.156) 0.313 (0.561) 1.47 (1.948) 14.871 (11.412)
$\theta_{2}$ 0.001 (0.002) 0.003 (0.006) 0.015 (0.019) 0.15 (0.112)
@@ E.2 Additional results for the Linear Regression case
We now present the results of the Linear Regression simulation in the well-specified case, i.e. when $\operatorname{\mathbb{F}}_{N}=\operatorname{\mathbb{F}}_{N}^{0}:=N(0,\sigma_{\nu}^{0})$ and $\operatorname{\mathbb{F}}_{X}=\operatorname{\mathbb{P}}_{X}^{0}:=N(0,1)$, for $\sigma_{\nu}^{0}\in\{10^{-6},0.5,1,2\}$. The linear regression fit is illustrated in Fig. 4 for all methods. As expected, the SIMEX method performs better than in the misspecified case whereas all other methods perform similarly.
@@
Robust-MEM $\theta_{1}$ 1.615 (0.129) 1.603 (0.166) 1.548 (0.273) 1.319 (0.573)
(TLS) $\theta_{2}$ 0.013 (0.002) 0.013 (0.003) 0.013 (0.004) 0.012 (0.006)
@@
E.3 Additional results for the Nonlinear Regression case with Berkson error
Similarly, we consider the example of nonlinear regression with Berkson ME from Section 5 in the well-specified case where both the true and the prior ME distributions are Gaussian with increasing variance. Results are reported for 100 simulations in Fig. 9.
@@ E.3 Additional results for the Nonlinear Regression case with Berkson error
We observe very similar behavior to the misspecified case. For ME variances of $\leq 0.25$, all methods exhibit comparable performance. However, as the variance escalates, the Robust-MEM (MMD) approach stands out as the most resilient to ME effects. This is also seen by observing the associated MSE and their variances in Table 7.
@@
Robust-MEM $\theta_{1}$ 0.139 (0.24) 0.127 (0.22) 0.13 (0.179) 0.244 (0.279)
(MMD) $\theta_{2}$ 0.442 (0.346) 0.238 (0.278) 0.4 (0.613) 1.264 (1.223)
@@
E.4 Results for Nonlinear Regression with Classical error
We now evaluate the nonlinear regression example for the Classical ME case. We report results for both the well-specified and misspecified cases of prior centering measures. In the well-specified case in (9), we set $\operatorname{\mathbb{F}}_{X}:=N(0,1)$ and $\operatorname{\mathbb{F}}_{W\>|\>X}:=N(x,\sigma^{2}_{\nu})$ where $\sigma^{2}_{\nu}$ is the true ME variance. The prior centering measure $\operatorname{\mathbb{F}}_{X\>|\>w_{i}}$ then has density $f_{X\>|\>w_{i}}(x\>|\>w_{i})\propto f_{W\>|\>X}(w_{i}\>|\>x)f_{X}(x)$ and is again a Normal distribution. In the misspecified case, we set $\operatorname{\mathbb{F}}_{X}:=N(0,1)$ and $\operatorname{\mathbb{F}}_{W\>|\>X}:=N(x,0.2)$. The obtained mean regression fit over 100 simulations is illustrated in Fig. 10 along with the estimates obtained from standard nonlinear least squares and the SIMEX procedure.
(a) ME variance is correctly specified
(b) ME variance is misspecified
Table 9: Mean Squared Error with associated standard deviation for $\theta=(\theta_{1},\theta_{2})$ in the nonlinear regression example with Classical ME of known ME variance (well-specified case). Results are reported for an increasing ME variance ($\sigma^{2}_{\nu}$) in the observed data, over 100 simulations.
Method ME variance
1e081e-08 $0.25$ $1.0$ $4.0$
LS $\theta_{1}$ 0.096 (0.179) 0.077 (0.093) 0.118 (0.201) 0.139 (0.127)
$\theta_{2}$ 0.264 (0.551) 0.385 (0.314) 1.455 (0.491) 2.926 (0.36)
R-MEM (MMD) $\theta_{1}$ 0.1 (0.117) 0.101 (0.118) 0.106 (0.122) 0.113 (0.128)
$\theta_{2}$ 0.059 (0.078) 0.056 (0.079) 0.058 (0.087) 0.06 (0.09)
SIMEX $\theta_{1}$ 0.096 (0.179) 0.111 (0.202) 0.261 (1.562) 0.13 (0.127)
$\theta_{2}$ 0.264 (0.552) 0.323 (0.663) 0.748 (1.019) 2.244 (0.576)
Table 10: Mean Squared Error with associated standard deviation for $\theta=(\theta_{1},\theta_{2})$ in the nonlinear regression example with Classical ME of unknown ME variance (misspecified case). Results are reported for an increasing ME variance ($\sigma^{2}_{\nu}$) in the observed data, over 100 simulations.
Method ME variance
1e081e-08 $0.25$ $1.0$ $4.0$
LS $\theta_{1}$ 0.096 (0.179) 0.077 (0.093) 0.118 (0.202) 0.139 (0.127)
$\theta_{2}$ 0.264 (0.551) 0.385 (0.314) 1.455 (0.491) 2.926 (0.36)
Robust-MEM $\theta_{1}$ 0.1 (0.12) 0.137 (0.154) 0.252 (0.236) 0.642 (0.408)
(MMD) $\theta_{2}$ 0.052 (0.063) 0.066 (0.08) 0.121 (0.134) 0.358 (0.295)
SIMEX $\theta_{1}$ 0.121 (0.27) 0.081 (0.093) 0.131 (0.31) 0.139 (0.128)
$\theta_{2}$ 0.46 (1.275) 0.337 (0.305) 1.414 (0.49) 2.919 (0.362)
ME variance
1e-08 0.5 1.0 2.0
Well-specified $\theta_{1}$ 0.96 0.97 0.98 0.98
$\theta_{2}$ 1 1 1 1
Misspecified $\theta_{1}$ 0.96 0.97 0.98 0.97
$\theta_{2}$ 1 1 1 1
E.5 Mental Health Study
In this section we give details of the model used in Harezlak et al., (2018) for the treatment effect estimation in the Mental Health study which we also adopt in Section 6.
For each observation $i=1,\dots,n_{placebo}$ in the placebo group we model
@@ E.5 Mental Health Study
We define the parameter of interest $\theta=(\beta,u_{0},u_{1})$ where
$\beta=(\beta_{0},\beta_{1},\beta_{0}^{\text{drug}},\beta_{1}^{\text{drug}}),u_{0}=(u_{01},\dots,u_{0K})\quad\text{and}\quad u_{1}=(u_{11},\dots,u_{1K}).$
Let $n=n_{treatment}+n_{placebo}$ and define $Y=(y_{1},\dots,y_{n})$ and the design matrices
$x=$ $\begin{pmatrix}1&x_{1}&1-I_{1}&(1-I_{1})x_{1}\\ \vdots&\vdots&\vdots&\vdots\\ 1&x_{n}&1-I_{n}&(1-I_{n})x_{n}\end{pmatrix},\quad Z_{0}=\begin{pmatrix}I_{1}z_{1}(x_{1})&\dots&I_{1}z_{K}(x_{1})\\ \vdots&\ddots&\vdots\\ I_{n}z_{1}(x_{n})&\dots&I_{n}z_{K}(x_{n})\end{pmatrix}$
@@
In this application we set $c=1$ and $T=100$ in the Robust-MEM algorithm. For the optimisation step we used a learning rate of $\eta=0.001$ and we employ a random restart method for the initialization of the optimisation: 100 points are sampled uniformly at random and the MMD loss is calculated using the observed data. The point with the smallest loss is used as the initial point and the resulting parameter estimate with the smallest loss during optimisation is picked. The uniform intervals are set to $[0,2]$ for the regression parameters $\beta$ and [0.1,0.1][-0.1,0.1] for the spline parameters $u_{0},u_{1}$. Posterior samples are obtained from 200 bootstrap iterations. Of course, when the optimisation objective is non-convex, there is usually a trade-off between coverage capability and MSE of the posterior estimates or intervals depending on the initialization of the optimisation. A wider uniform interval usually leads to higher uncertainty in our posterior beliefs which can exhibit better coverage probability of estimated intervals. On the contrary, if one is interested in the MSE of the posterior estimations then one can perform the above procedure more than 100 times to get a better initial point of the optimisation or use a narrower uniform distribution based on prior beliefs or historic data.